A battery operated car of mass 1,236 kilograms uses a 17.3 volt system. What is the charge in coulombs the batteries must be able to move in order to make the car climb a hill that is 626 meters high?
Answer #1

energy required = mgh

= 1236*9.8*626 = 7582612.8 J

7582612.8 = VIt

7582612.8 = V*q

q = 7582612.8 / 17.3 = 4.38*10^5 C

so the charge = 4.38*10^5 C